### v3

@@ 5.1 Sparse Spectral Clustering (SSC)
By Corollary 5.1, the problem (77) can be reformulated as Problem 1.2 with $C\coloneqq{\rm St}(K,N)$, $\bm{\varphi}\coloneqq\bm{\varphi}_{\bm{S}}^{\rm Cay}$ in (72), $h(\bm{U})\coloneqq{\rm Tr}(\bm{U}^{\mathsf{T}}\bm{L}\bm{U})$ and $g\coloneqq\lambda\psi$:
@@
In this numerical experiment, we evaluated the numerical performance of SSC via the formulation in (78) by applying Algorithm 1 with $\gamma_{n}$ given by Algorithm 2 under the setting (for a choice of $(\mu_{n})_{n=1}^{\infty}$, see Remark 4.10)
$(\mu_{n})_{n=1}^{\infty}=((2\eta)^{-1}n^{-1/3})_{n=1}^{\infty},c=2^{-13},\rho=0.5,\gamma_{\rm initial}=\min\left\{1,\left\lVert\nabla(F_{1}\circ\bm{\varphi}_{\bm{S}}^{\rm Cay})(\bm{V}_{1})\right\rVert_{F}^{-1}\right\},$ (79)
where $\left\lVert\cdot\right\rVert_{F}$ is the Frobenius norm, $F_{1}$ is defined as in Algorithm 1, and $(\bm{S},\bm{V}_{1})\in{\rm O}(N)\times Q_{N,K}$ was chosen [^15] By following [54, Thm. 2.7], a pair $(\bm{S},\bm{V}_{1})\in{\rm O}(N)\times Q_{N,K}$ satisfying $\bm{\varphi}_{\bm{S}}^{\rm Cay}(\bm{V}_{1})=\bm{U}_{1}$ can be chosen as $\bm{S}\coloneqq\mathrm{diag}(\bm{Q}_{1}\bm{Q}_{2}^{\mathsf{T}},\bm{I}_{N-K})$ and $\bm{V}_{1}\coloneqq\Phi_{\bm{S}}(\bm{U}_{1})$, where $\bm{Q}_{1},\bm{Q}_{2}\in{\rm O}(K)$ are obtained by a singular-value decomposition of the upper block matrix $\bm{U}_{0{\rm up}}=\bm{Q}_{1}\bm{\Sigma}\bm{Q}_{2}\in\mathbb{R}^{K\times K}$ of $\bm{U}_{1}$ with a nonnegative-valued diagonal matrix $\bm{\Sigma}\in\mathbb{R}^{K\times K}$, and $\Phi_{\bm{S}}:{\rm St}(K,N)\setminus E_{N,K}(\bm{S})\to Q_{N,K}$ is the generalized Cayley transform [54, (10)-(12)] (Note: $\Phi_{\bm{S}}\circ\bm{\varphi}_{\bm{S}}^{\rm Cay}=\mathrm{Id}$ and $\bm{\varphi}_{\bm{S}}^{\rm Cay}\circ\Phi_{\bm{S}}=\mathrm{Id}$ hold respectively on their domains). to satisfy $\bm{\varphi}_{\bm{S}}^{\rm Cay}(\bm{V}_{1})=\bm{U}_{1}$ for an initial guess $\bm{U}_{1}\in{\rm St}(K,N)$. In order to examine how much weakly convex regularizers can enhance the performance, we compared (i) the proposed SSC (SSC+MCP) by employing MCP [16]
@@ 5.2 Sparse Principal Component Analysis (SPCA)
where $\bm{\Xi}\in\mathbb{R}^{\mathcal{I}\times N}$ is a data matrix, $\mathcal{I}\in\mathbb{N}$ is the number of data, $N$ is the dimension of each data, and $\lambda>0$ is a predetermined weight. The SPCA in (83) is a special instance of Problem 1.1 with $C\coloneqq{\rm St}(p,N)$ and $\mathfrak{S}=\mathrm{Id}$. By Corollary 5.1, the SPCA in (83) can be reformulated as Problem 1.2 with $C\coloneqq{\rm St}(p,N)$ and $\bm{\varphi}\coloneqq\bm{\varphi}_{\bm{S}}^{\rm Cay}$ in (72):
@@
We compared the numerical performance of VSmooth with that of three Riemannian nonsmooth optimization algorithms applied to the problem in (83): (i) the Riemannian smoothing gradient algorithm [24, Section 4.1] (RSmooth), (ii) the Riemannian subgradient method [21, (2.6)] (RSub), and (iii) the Riemannian proximal gradient method [20, Algorithm 2] (ManPGAda [^18] We used codes in. https://github.com/chenshixiang/ManPG. ). For each Riemannian optimization algorithm, the update from $\bm{U}_{n}\in{\rm St}(p,N)$ to $\bm{U}_{n+1}\in{\rm St}(p,N)$ is expressed as
@@
where $P_{T_{{\rm St}(p,N)}(\bm{U}_{n})}:\mathbb{R}^{N\times p}\to T_{{\rm St}(p,N)}(\bm{U}_{n})$ is the orthogonal projection mapping onto $T_{{\rm St}(p,N)}(\bm{U}_{n})$ (see, e.g., [30]), and $t_{n}>0$ for ManPGAda [20] is adaptively designed to accelerate convergence speed. Clearly, ManPGAda requires some iterative solver for the subproblem (89) to find a search direction while the others do not require, i.e., the others are single-loop algorithms. For RSmooth, we employed $(\mu_{n})_{n=1}^{\infty}=((2\eta)^{-1}n^{-1/3})_{n=1}^{\infty}$ because RSmooth with this $(\mu_{n})_{n=1}^{\infty}$ achieves the best convergence rate [24, Remark 4.1]. In this experiment, to find a stepsize $\gamma_{n}$, we used backtracking algorithms for ManPGAda and RSmooth suggested in [20] and [24] respectively, and we employed $\gamma_{n}\coloneqq 0.99^{n}$ for RSub.
@@
From Figure 5.2, we see that the proposed VSmooth converges faster than RSmooth and RSub for all problem sizes. Thus, VSmooth achieves the best performance among algorithms that do not require any iterative solver for subproblems. For $p\in\{1,N/100\}$, VSmooth seems to be competitive or inferior to ManPGAda. However, for $N\in\{500,1000\}$ and large $p=N/10$, the performance of ManPGAda severely deteriorates while no such deterioration is observed for VSmooth. Moreover, from Table 4, we also observe that ManPGAda updates for only a few iterations; and CPU time of ManPGAda significantly exceeds the time limitation shown by Table 3. These observations imply that ManPGAda took a long time to find an approximated solution of a subproblem in (89) at every iteration, and thus single-loop algorithms, namely, VSmooth, RSmooth and RSub, are more computationally reliable than ManPGAda for large problem sizes. From this experimental result, VSmooth has a comparable numerical performance compared to the fairly standard Riemannian nonsmooth optimization algorithms [20, 21, 24] even for Problem 1.1 with a simple case, i.e., $g$ is convex and $\mathfrak{S}=\mathrm{Id}$, where the proposed VSmooth can be applied to Problem 1.1, via Problem 1.2, with more general setting of the nonlinearity of $\mathfrak{S}$ as seen in Section 5.1.
@@ 6 Conclusions
We addressed a manifold constrained nonsmooth optimization problem, where the cost function consists of a smooth function and the composite function of a weakly convex function and a smooth (nonlinear) mapping. For the target problem, we proposed a variable smoothing algorithm with a parametrization of a nonconvex constraint set. We also presented an asymptotic convergence analysis and a convergence rate analysis of the proposed algorithm in terms of stationary points. The numerical experiments demonstrate that (i) the proposed algorithm improves the numerical performance of sparsity-aware application by using a weakly convex regularizer; (ii) the proposed algorithm has a comparable numerical performance compared with the existing Riemannian nonsmooth optimization algorithms based on a retraction.
@@
Appendix A Proof of Lemma 2.4
@@
Although Lemma 2.4 can be checked by the subdifferential calculus rules [57, Thm. 10.6, Cor. 10.9], we present a short proof of Lemma 2.4 with some necessary facts in order to increase the readability for researchers in the area of applications. The subdifferential calculus rules heavily rely on the horizontal subdifferential, which gives constraint qualifications, and the subdifferential regularity of a function (see Definition 2.2 (b) and Example A.2).
Definition A.1 (Horizontal subdifferential [57, Def. 8.3 and Thm. 8.9]).
@@
Example A.2 (Subdifferentially regular functions).
Subdifferentially regular functions include, e.g., continuously differentiable functions [57, Exm. 7.28], proper lower semicontinuous (weakly) convex functions [57, Exm. 7.27] [86, Prop. 4.4.15], and indicator functions of Clarke regular subsets [57, Exm. 7.28].
Fact A.3 presents useful properties, regarding $\partial^{\infty}J$, that will be used for our analyses.
Fact A.3 (Horizontal subdifferential being zero [57, Thm. 9.13]).
@@
Fact A.4 (Calculus rules for subdifferential).
@@
    (Chain rule [57, Thm. 10.6]) If the constraint qualification $\partial^{\infty}J_{1}(\mathcal{F}(\widebar{\bm{y}}))\cap\mathrm{Ker}((\mathrm{D}\mathcal{F}(\widebar{\bm{y}}))^{*})=\{\bm{0}\}$ holds at $\widebar{\bm{y}}\in\mathrm{dom}(J_{1}\circ\mathcal{F})$, then $J_{1}\circ\mathcal{F}$ is subdifferentially regular at $\widebar{\bm{y}}$ and
@@
    (Sum rule [57, Cor. 10.9, Exe. 8.8]) If the constraint qualification $\partial^{\infty}J_{1}(\widebar{\bm{x}})\cap(-\partial^{\infty}J_{2}(\widebar{\bm{x}}))=\{\bm{0}\}$ holds at $\widebar{\bm{x}}\in\mathrm{dom}(J_{1}+J_{2})$, then $J_{1}+J_{2}$ is subdifferentially regular at $\widebar{\bm{x}}$, and
@@
(Step 2) Recall that $\iota_{C}$ is subdifferentially regular (see Example A.2 and Example 2.3), and lower semicontinuous due to the closedness of $C$. Then, by letting $J_{1}\coloneqq F$, $J_{2}\coloneqq\iota_{C}$, and $\mathcal{F}\coloneqq\bm{\varphi}$, Fact A.4 yields that (i) the subdifferential regularities of $F\circ\bm{\varphi}$ and $F+\iota_{C}$; (ii) the equalities (17) and (18) from (92) and (90), where $J_{1}$ is subdifferentially regular and $\partial^{\infty}J_{1}(\widebar{\bm{x}})=\{\bm{0}\}\ (\widebar{\bm{x}}\in\mathcal{X})$ from (Step 1). ∎
Appendix B Proof of Proposition 4.2
@@
Lemma B.1 (Lipschitz continuity of gradient of composite function).
@@
    The operator norm of $(\mathrm{D}\mathcal{F}(\cdot))^{*}$ is bounded above by $\kappa_{\mathcal{F}}>0$ over $\mathrm{Conv}(E)$, and thus $\mathcal{F}$ is Lipschitz continuous with a Lipschitz constant $\kappa_{\mathcal{F}}$ over $\mathrm{Conv}(E)$ by [57, Thm. 9.2 and Thm. 9.7].
@@
(a) Since $\nabla{}^{\mu}g$ is Lipschitz continuous with the Lipschitz constant $L_{\nabla{}^{\mu}g}$ and $\left\lVert\nabla{}^{\mu}g(\bm{z})\right\rVert\leq L_{g}\ (\bm{z}\in\mathcal{Z})$ by (22), Lemma B.1 with setting $(\mathcal{H},\mathcal{K},\mathcal{F},J,E)\coloneqq(\mathcal{Z},\mathcal{X},\mathfrak{S},{}^{\mu}g,C)$ ensures the Lipschitz continuity of $\nabla({}^{\mu}g\circ\mathfrak{S})$ with a Lipschitz constant $L_{\nabla({}^{\mu}g\circ\mathfrak{S})}\coloneqq\kappa_{\mathfrak{S}}^{2}L_{\nabla{}^{\mu}g}+L_{g}L_{\mathrm{D}\mathfrak{S}}=\kappa_{\mathfrak{S}}^{2}\mu^{-1}+L_{g}L_{\mathrm{D}\mathfrak{S}}$. Thus, $\nabla(h+{}^{\mu}g\circ\mathfrak{S})$ is Lipschitz continuous over $C$ with a Lipschitz constant $L_{\nabla(h+{}^{\mu}g\circ\mathfrak{S})}=L_{\nabla h}+L_{g}L_{\mathrm{D}\mathfrak{S}}+\kappa_{\mathfrak{S}}^{2}\mu^{-1}$, where $L_{\nabla h}$ is the Lipschitz constant of $\nabla h$ by the setting of Problem 1.2.
(b) To apply Lemma B.1, we derive an upper bound $\kappa_{h+{}^{\mu}g\circ\mathfrak{S}}\coloneqq\kappa_{h}+\kappa_{\mathfrak{S}}L_{g}>0$ of $\sup_{\bm{x}\in C}\left\lVert\nabla(h+{}^{\mu}g\circ\mathfrak{S})(\bm{x})\right\rVert$ as:
@@
Appendix C Proof of Theorem 4.4
@@
Lemma C.1 (Subdifferential limit).

### v4

@@ 5.1 Sparse Spectral Clustering (SSC)
By Corollary 5.1, the problem (82) can be reformulated as Problem 1.4 with $C\coloneqq{\rm St}(K,N)$, $\bm{\varphi}\coloneqq\bm{\varphi}_{\bm{S}}^{\rm Cay}$ in (77), $h(\bm{U})\coloneqq{\rm Tr}(\bm{U}^{\mathsf{T}}\bm{L}\bm{U})$ and $g\coloneqq\lambda\psi$:
@@
In this numerical experiment, we evaluated the numerical performance of SSC via the formulation in (83) by applying Algorithm 1 with $\gamma_{n}$ given by Algorithm 2 under the setting (for a choice of $(\mu_{n})_{n=1}^{\infty}$, we follow the setting in the oracle complexity result; see Corollary 4.12)
$(\mu_{n})_{n=1}^{\infty}=((2\eta)^{-1}n^{-1/3})_{n=1}^{\infty},c=2^{-13},\rho=0.5,\gamma_{\rm init}=\min\left\{1,\left\lVert\nabla(F_{1}\circ\bm{\varphi}_{\bm{S}}^{\rm Cay})(\bm{V}_{1})\right\rVert_{F}^{-1}\right\},$ (84)
where $\left\lVert\cdot\right\rVert_{F}$ is the Frobenius norm, $F_{1}$ is defined as in Algorithm 1, and $(\bm{S},\bm{V}_{1})\in{\rm O}(N)\times Q_{N,K}$ was chosen [^17] By following [48, Thm. 2.7], a pair $(\bm{S},\bm{V}_{1})\in{\rm O}(N)\times Q_{N,K}$ satisfying $\bm{\varphi}_{\bm{S}}^{\rm Cay}(\bm{V}_{1})=\bm{U}_{1}$ can be chosen as $\bm{S}\coloneqq\mathrm{diag}(\bm{Q}_{1}\bm{Q}_{2}^{\mathsf{T}},\bm{I}_{N-K})$ and $\bm{V}_{1}\coloneqq\Phi_{\bm{S}}(\bm{U}_{1})$, where $\bm{Q}_{1},\bm{Q}_{2}\in{\rm O}(K)$ are obtained by a singular-value decomposition of the upper block matrix $\bm{U}_{0{\rm up}}=\bm{Q}_{1}\bm{\Sigma}\bm{Q}_{2}\in\mathbb{R}^{K\times K}$ of $\bm{U}_{1}$ with a nonnegative-valued diagonal matrix $\bm{\Sigma}\in\mathbb{R}^{K\times K}$, and $\Phi_{\bm{S}}:{\rm St}(K,N)\setminus E_{N,K}(\bm{S})\to Q_{N,K}$ is the generalized Cayley transform [48, (10)-(12)] (Note: $\Phi_{\bm{S}}\circ\bm{\varphi}_{\bm{S}}^{\rm Cay}=\mathrm{Id}$ and $\bm{\varphi}_{\bm{S}}^{\rm Cay}\circ\Phi_{\bm{S}}=\mathrm{Id}$ hold respectively on their domains). to satisfy $\bm{\varphi}_{\bm{S}}^{\rm Cay}(\bm{V}_{1})=\bm{U}_{1}$ for an initial guess $\bm{U}_{1}\in{\rm St}(K,N)$. In order to examine how much weakly convex regularizers can enhance the performance, we compared (i) the proposed SSC (SSC+MCP) by employing MCP [89]
@@ 5.2 Sparse Principal Component Analysis (SPCA)
where $\bm{\Xi}\in\mathbb{R}^{\mathcal{I}\times N}$ is a data matrix, $\mathcal{I}\in\mathbb{N}$ is the number of data, $N$ is the dimension of each data, and $\lambda>0$ is a predetermined weight. The SPCA in (88) is a special instance of Problem 1.1 with $C\coloneqq{\rm St}(p,N)$ and $\mathfrak{S}=\mathrm{Id}$. By Corollary 5.1, the SPCA in (88) can be reformulated as Problem 1.4 with $C\coloneqq{\rm St}(p,N)$ and $\bm{\varphi}\coloneqq\bm{\varphi}_{\bm{S}}^{\rm Cay}$ in (77):
@@
We compared the numerical performance of VSmooth with that of three Riemannian nonsmooth optimization algorithms applied to the problem in (88): (i) the Riemannian smoothing gradient algorithm [10, Section 4.1] (RSmooth), (ii) the Riemannian subgradient method [59, (2.6)] (RSub), and (iii) the Riemannian proximal gradient method [25, Algorithm 2] (ManPGAda [^20] We used codes in. https://github.com/chenshixiang/ManPG. ). For each Riemannian optimization algorithm, the update from $\bm{U}_{n}\in{\rm St}(p,N)$ to $\bm{U}_{n+1}\in{\rm St}(p,N)$ is expressed as
@@
where $P_{T_{{\rm St}(p,N)}(\bm{U}_{n})}:\mathbb{R}^{N\times p}\to T_{{\rm St}(p,N)}(\bm{U}_{n})$ is the orthogonal projection mapping onto $T_{{\rm St}(p,N)}(\bm{U}_{n})$ (see, e.g., [3]), and $t_{n}>0$ for ManPGAda [25] is adaptively designed to accelerate convergence speed. Clearly, ManPGAda requires some iterative solver for the subproblem (94) to find a search direction while the others do not require, i.e., the others are single-loop algorithms. For RSmooth, we employed $(\mu_{n})_{n=1}^{\infty}=((2\eta)^{-1}n^{-1/3})_{n=1}^{\infty}$ because RSmooth with this $(\mu_{n})_{n=1}^{\infty}$ achieves the best convergence rate [10, Remark 4.1]. In this experiment, to find a stepsize $\gamma_{n}$, we used backtracking algorithms for ManPGAda and RSmooth suggested in [25] and [10] respectively, and we employed $\gamma_{n}\coloneqq 0.99^{n}$ for RSub.
@@
From Figure 5.2, we see that the proposed VSmooth converges faster than RSmooth and RSub for all problem sizes. Thus, VSmooth achieves the best performance among algorithms that do not require any iterative solver for subproblems. For $p\in\{1,N/100\}$, VSmooth seems to be competitive or inferior to ManPGAda. However, for $N\in\{500,1000\}$ and large $p=N/10$, the performance of ManPGAda severely deteriorates while no such deterioration is observed for VSmooth. Moreover, from Table 4, we also observe that ManPGAda updates for only a few iterations; and CPU time of ManPGAda significantly exceeds the time limitation shown by Table 3. These observations imply that ManPGAda took a long time to find an approximated solution of a subproblem in (94) at every iteration, and thus single-loop algorithms, namely, VSmooth, RSmooth and RSub, are more computationally reliable than ManPGAda for large problem sizes. From this experimental result, VSmooth has a comparable numerical performance compared to the fairly standard Riemannian nonsmooth optimization algorithms [25, 59, 10] even for Problem 1.1 with a simple case, i.e., $g$ is convex and $\mathfrak{S}=\mathrm{Id}$, where the proposed VSmooth can be applied to Problem 1.1, via Problem 1.4, with more general setting of the nonlinearity of $\mathfrak{S}$ as seen in Section 5.1.
@@ 6 Conclusions
We addressed a manifold constrained nonsmooth optimization problem, where the cost function consists of a smooth function and the composite function of a weakly convex function and a smooth (nonlinear) mapping. For the target problem, we proposed a variable smoothing algorithm with a parametrization of a nonconvex constraint set. We also presented an asymptotic convergence analysis and a first-order oracle complexity $\mathcal{O}(\epsilon^{-3})$ of the proposed algorithm in terms of stationary points. The numerical experiments demonstrate that (i) the proposed algorithm improves the numerical performance of sparsity-aware application by using a weakly convex regularizer; (ii) the proposed algorithm has a comparable numerical performance compared with the existing Riemannian nonsmooth optimization algorithms based on a retraction.
@@
Appendix A Related Work
A.1 Smoothing-type algorithms
We review smoothing-type algorithms for finding a stationary point of a nonsmooth function $J:\mathcal{X}\to\mathbb{R}$ because the proposed algorithm (Algorithm 1) can be viewed as a variant of smoothing-type algorithms. Smoothing methods have been extensively developed [87, 26, 44, 79, 53, 88, 63] with a smoothed surrogate function, say $J^{\langle\mu\rangle}:\mathcal{X}\to\mathbb{R}\ (\mu\in\mathbb{R}_{++})$, having the (sub-)gradient consistency property (see Remark 4.5) with $\lim_{\mathbb{R}_{++}\ni\mu\to 0}J^{\langle\mu\rangle}(\bm{x})=J(\bm{x})\ (\bm{x}\in\mathcal{X})$. The gradient consistency property has been utilized as a key tool to express the first-order optimality condition in terms of a gradient sequence of $\nabla J^{\langle\mu\rangle}$ with $\mu\searrow 0$. The basic update of a smoothing gradient method [26] is given by $\bm{x}_{n+1}\coloneqq\bm{x}_{n}-\gamma_{n}\nabla J^{\langle\mu_{n}\rangle}(\bm{x}_{n})\ (n\in\mathbb{N})$ with a properly chosen stepsize $\gamma_{n}\in\mathbb{R}_{++}$, where smoothing parameters $\mu_{n}$ are conditionally updated with $\sigma\in(0,1)$ and $\delta\in\mathbb{R}_{++}$ by setting $\mu_{n+1}\coloneqq\sigma\mu_{n}$ if $\left\lVert\nabla J^{\langle\mu_{n}\rangle}(\bm{x}_{n})\right\rVert<\delta\mu_{n}$; $\mu_{n+1}\coloneqq\mu_{n}$ otherwise. Thanks to the (sub-)gradient consistency property of $J^{\langle\mu\rangle}$, a cluster point of $(\bm{x}_{n})_{n=1}^{\infty}$ is guaranteed to be a (Clarke) stationary point of $J$ [26]. Based on the conditional update of smoothing parameters, many numerical algorithms, e.g., accelerated gradient methods [79, 63], projected gradient methods [87, 44] and manifold constrained smoothing methods [53, 88], for updating $(\bm{x}_{n})_{n=1}^{\infty}$ have been established with asymptotic convergence analyses in terms of a stationary point. However, these smoothing methods can be viewed as double-loop algorithms due to the conditional update of smoothing parameters.
Another class of smoothing-type methods has been proposed with iteratively decreasing smoothing parameters $(\mu_{n})_{n=1}^{\infty}$ [15, 74, 14, 13, 46, 81, 10, 66, 64], where these methods are called variable/adaptive/dynamic smoothing methods. Thanks to the use of iteratively decreasing smoothing parameters, these variable smoothing methods are single-loop algorithms. For convex optimization problems, the oracle complexities for finding the optimal value of the cost function have been shown for [15, 74, 14, 46, 81, 64], where an exceptional paper [81] guarantees that a generated sequence converges to a minimizer of the cost function. Recently, for nonconvex nonsmooth optimization problems, a variable smoothing algorithm [15] has been extended by [13] to a special version of Problem 1.1 where $\mathfrak{S}$ is a linear operator and $C\coloneqq\mathcal{X}$. The variable smoothing algorithm in [13] uses a smoothed surrogate function $h+{}^{\mu}g\circ\mathfrak{S}$ with the Moreau envelope $^{\mu}g$ of $g$ (see (21)), and its first-order oracle complexity $\mathcal{O}(\epsilon^{-3})$ for finding a certain $\epsilon$-approximate stationary point in the sense of (73) has been shown under the linearity of $\mathfrak{S}$. This variable smoothing algorithm [13] with an oracle complexity guarantee has also been extended by [10, 66] to Problem 1.1 where $C$ is a compact embedded submanifold of $\mathcal{X}$, but $\mathfrak{S}$ is still assumed to be linear. However, for variable smoothing algorithms [13, 66], except for [10], any asymptotic convergence analysis to a stationary point has not been reported yet, where [10] presents an asymptotic convergence analysis by exploiting the compactness of $C$ and the linearity of $\mathfrak{S}$ (see also Table 1 in Section 1).
A.2 Penalty-based strategy
In a case where $C$ in Problem 1.1 is defined by equality constraints, i.e., $C=H^{-1}(\bm{0})$ with some continuously differentiable mapping $H:\mathcal{X}\to\mathbb{R}^{m}$, penalty-based strategies have been proposed, e.g., [65, Ch. 12][82], as an alternative strategy to the parametrization strategy in Section 3 for translating Problem 1.1 into an unconstrained problem. Such penalty-based strategies aim to find a point satisfying the Karush-Kuhn-Tucker (KKT) condition by assuming the availability of $H$ with suitable regularities, e.g., Linearly Independent Constraint Qualification (LICQ), i.e., $\mathrm{rank}(\mathrm{D}H(\bm{x}))=m$ for all $\bm{x}\in C$, where $H$ with LICQ is called a (global) defining map [16, Cor. 8.76]. Indeed, LICQ of $H$ makes the KKT condition a necessary condition for the local optimality of Problem 1.1 with a smooth $F$ [65, Thm. 12.1], while the KKT condition without LICQ of $H$ is not guaranteed to be a necessary condition for the local optimality. However, the existence of a global defining map $H$ with LICQ is not guaranteed for general embedded submanifolds [16, p.50], e.g., non-orientable manifolds [54, Prop. 15.23, pp.106-107] (see [76] for applications of optimization over non-orientable manifolds). Moreover, penalty-based strategies [65, 82] may generate an infeasible sequence $(\bm{x}_{n})_{n=1}^{\infty}\subset\mathcal{X}$, i.e., $\bm{x}_{n}$ may not belong to $C$, where the feasibility is guaranteed at cluster points of $(\bm{x}_{n})_{n=1}^{\infty}$ in general.
In contrast, for an embedded submanifold $C$ in Problem 1.1, a parametrization $\bm{\varphi}$ of $C$ exists (see Example 1.3 (c)), and a stationary point $\bm{y}^{\star}\in\mathcal{Y}$ of Problem 1.4 satisfies a necessary condition for the local optimality of Problem 1.1 at $\bm{x}^{\star}\coloneqq\bm{\varphi}(\bm{y}^{\star})\in C$ (see Remark 3.6). Under an additional condition, e.g., submersion condition on $\bm{\varphi}$, $\bm{x}^{\star}$ is a stationary point of Problem 1.1 (see Theorem 3.2 and Corollary 3.4). Furthermore, the parametrization strategy described in Section 3 generates a sequence $(\bm{y}_{n})_{n=1}^{\infty}\subset\mathcal{Y}$ for finding a stationary point $\bm{y}^{\star}$; hence the feasibility of $(\bm{\varphi}(\bm{y}_{n}))_{n=1}^{\infty}\subset C$ is guaranteed via $\bm{\varphi}$ (see Algorithm 1 in Section 4.2; see also, e.g., [84, 36, 43, 41, 58, 47, 48, 56], for parametrization strategy with smooth cost function).
Finally, we remark that the submersion condition of a parametrization $\bm{\varphi}$ and LICQ of $H$ are closely related if both $\bm{\varphi}$ with the submersion condition and a global defining map $H$ with LICQ exist for an embedded submanifold $C$. Indeed, in this case, we have $T_{C}(\bm{x})=\mathrm{Ker}(\mathrm{D}H(\bm{x}))=\mathrm{Ran}(\mathrm{D}\bm{\varphi}(\bm{y}))\ (\bm{y}\in\mathcal{Y},\bm{x}\coloneqq\bm{\varphi}(\bm{y})\in C)$. However, the availabilities of $\bm{\varphi}$ with the submersion condition and a global defining map $H$ with LICQ are not comparable in general because we have an example [^21] The Klein bottle has no global defining map satisfying LICQ due to its non-orientability (see, e.g., [76, Sect. 3.3.1] and [54, Prop. 15.23, pp.106-107]). However, since $\mathbb{R}^{2}$ is the universal covering manifold [54, p.90] of the Klein bottle (see, e.g., [6]), there exists a smooth covering mapping [54, p.90] from $\mathbb{R}^{2}$ onto the Klein bottle. Such a smooth covering mapping is a parametrization $\bm{\varphi}$, defined over $\mathbb{R}^{2}$, of the Klein bottle such that the submersion condition is satisfied at every point in $\mathbb{R}^{2}$ by [54, Prop. 4.33 (a)]. of $C$ such that any global defining map $H$ of $C$ with LICQ does not exist, whereas a parametrization $\bm{\varphi}:\mathcal{Y}\to\mathcal{X}$ of $C$ with the submersion condition exists.
Appendix B Proof of Lemma 2.4
@@ A.2 Penalty-based strategy
Although Lemma 2.4 can be checked by the subdifferential calculus rules [69, Thm. 10.6, Cor. 10.9], we present a short proof of Lemma 2.4 with some necessary facts in order to increase the readability for researchers in the area of applications. The subdifferential calculus rules heavily rely on the horizontal subdifferential, which gives constraint qualifications, and the subdifferential regularity of a function (see Definition 2.2 (b) and Example B.2).
Definition B.1 (Horizontal subdifferential [69, Def. 8.3 and Thm. 8.9]).
@@
Example B.2 (Subdifferentially regular functions).
Subdifferentially regular functions include, e.g., continuously differentiable functions [69, Exm. 7.28], proper lower semicontinuous (weakly) convex functions [69, Exm. 7.27] [30, Prop. 4.4.15], and indicator functions of Clarke regular subsets [69, Exm. 7.28].
Fact B.3 presents useful properties, regarding $\partial^{\infty}J$, that will be used for our analyses.
Fact B.3 (Horizontal subdifferential being zero [69, Thm. 9.13]).
@@
Fact B.4 (Calculus rules for subdifferential).
@@
    (Chain rule [69, Thm. 10.6]) If the constraint qualification $\partial^{\infty}J_{1}(\mathcal{F}(\widebar{\bm{y}}))\cap\mathrm{Ker}((\mathrm{D}\mathcal{F}(\widebar{\bm{y}}))^{*})=\{\bm{0}\}$ holds at $\widebar{\bm{y}}\in\mathrm{dom}(J_{1}\circ\mathcal{F})$, then $J_{1}\circ\mathcal{F}$ is subdifferentially regular at $\widebar{\bm{y}}$ and
@@
    (Sum rule [69, Cor. 10.9, Exe. 8.8]) If the constraint qualification $\partial^{\infty}J_{1}(\widebar{\bm{x}})\cap(-\partial^{\infty}J_{2}(\widebar{\bm{x}}))=\{\bm{0}\}$ holds at $\widebar{\bm{x}}\in\mathrm{dom}(J_{1}+J_{2})$, then $J_{1}+J_{2}$ is subdifferentially regular at $\widebar{\bm{x}}$, and
@@
(Step 2) Recall that $\iota_{C}$ is subdifferentially regular (see Example B.2 and Example 2.3), and lower semicontinuous due to the closedness of $C$. Then, by letting $J_{1}\coloneqq F$, $J_{2}\coloneqq\iota_{C}$, and $\mathcal{F}\coloneqq\bm{\varphi}$, Fact B.4 yields that (i) the subdifferential regularities of $F\circ\bm{\varphi}$ and $F+\iota_{C}$; (ii) the equalities (17) and (18) from (97) and (95), where $J_{1}$ is subdifferentially regular and $\partial^{\infty}J_{1}(\widebar{\bm{x}})=\{\bm{0}\}\ (\widebar{\bm{x}}\in\mathcal{X})$ from (Step 1). ∎
Appendix C Proof of Proposition 4.2
@@
Lemma C.1 (Lipschitz continuity of gradient of composite function).
@@
    The operator norm of $(\mathrm{D}\mathcal{F}(\cdot))^{*}$ is bounded above by $\kappa_{\mathcal{F}}>0$ over $\mathrm{Conv}(E)$, and thus $\mathcal{F}$ is Lipschitz continuous with a Lipschitz constant $\kappa_{\mathcal{F}}$ over $\mathrm{Conv}(E)$ by [69, Thm. 9.2 and Thm. 9.7].
@@
(a) Since $\nabla{}^{\mu}g$ is Lipschitz continuous with the Lipschitz constant $L_{\nabla{}^{\mu}g}$ and $\left\lVert\nabla{}^{\mu}g(\bm{z})\right\rVert\leq L_{g}\ (\bm{z}\in\mathcal{Z})$ by (22), Lemma C.1 with setting $(\mathcal{H},\mathcal{K},\mathcal{F},J,E)\coloneqq(\mathcal{Z},\mathcal{X},\mathfrak{S},{}^{\mu}g,C)$ ensures the Lipschitz continuity of $\nabla({}^{\mu}g\circ\mathfrak{S})$ with a Lipschitz constant $L_{\nabla({}^{\mu}g\circ\mathfrak{S})}\coloneqq\kappa_{\mathfrak{S}}^{2}L_{\nabla{}^{\mu}g}+L_{g}L_{\mathrm{D}\mathfrak{S}}=\kappa_{\mathfrak{S}}^{2}\mu^{-1}+L_{g}L_{\mathrm{D}\mathfrak{S}}$. Thus, $\nabla(h+{}^{\mu}g\circ\mathfrak{S})$ is Lipschitz continuous over $C$ with a Lipschitz constant $L_{\nabla(h+{}^{\mu}g\circ\mathfrak{S})}=L_{\nabla h}+L_{g}L_{\mathrm{D}\mathfrak{S}}+\kappa_{\mathfrak{S}}^{2}\mu^{-1}$, where $L_{\nabla h}$ is the Lipschitz constant of $\nabla h$ by the setting of Problem 1.4.
(b) To apply Lemma C.1, we derive an upper bound $\kappa_{h+{}^{\mu}g\circ\mathfrak{S}}\coloneqq\kappa_{h}+\kappa_{\mathfrak{S}}L_{g}>0$ of $\sup_{\bm{x}\in C}\left\lVert\nabla(h+{}^{\mu}g\circ\mathfrak{S})(\bm{x})\right\rVert$ as:
@@
Appendix D Proof of Theorem 4.4
@@
Lemma D.1 (Subdifferential limit).